# Event Topology-based Visual Microphone for Amplitude and Frequency Reconstruction

### v1

@@ Abstract
Event cameras are also particularly well suited for vibration measurement owing to their ability to capture fast dynamics. Approaches for vibration reconstruction using an event camera can broadly be categorized as active or passive capture. Active methods, which project an external light source onto a target, can realize high-sensitivity recordings. However, their use of lasers introduces limitations Yin et al. (2025); Cai et al. (2025). Although this approach is precise for single-point measurements, its reliance on capturing laser speckle patterns introduces practical limitations. For instance, the need to defocus the laser makes it difficult to record multiple vibration sources simultaneously, preventing the full utilization of the event camera’s high spatial resolution. Thus, we focus on passive methods, which record the target without using a strong light source. One approach to vibration reconstruction is interpreting the times of specular reflections in event data as signal level crossings and recovering the signal through short-time Fourier sparsity Howard and Hirakawa (2023); Howard, Jones, and Hirakawa (2025). In our previous work, we also reconstructed vibrations by generating pseudo-frames from the event data and extracting phase information using Gabor filters Niwa et al. (2023). Although these passive methods can successfully reconstruct the principal frequency components of a vibration, they fail to accurately recover vibration amplitudes.
@@
Next, a clustering algorithm is applied to the subset of events within each colored interval. This process partitions the events into distinct clusters. For this task, we employ HDBSCAN (Hierarchical Density-Based Spatial Clustering of Applications with Noise) Campello, Moulavi, and Sander (2013). Unlike conventional density-based clustering such as DBSCAN Ester et al. (1996), which requires a manually tuned distance threshold $\varepsilon$, HDBSCAN adapts to varying event densities. This adaptability is crucial for robustly handling complex vibrations where frequency and amplitude changes alter the local event density. Figure 2(c) illustrates the outcome of this stage, showing the identified clusters after HDBSCAN has filtered out sparse events classified as noise. The interval and HDBSCAN parameters are detailed in the supplementary material.
@@
To evaluate the reconstruction accuracy, we employ two primary metrics. Because the reconstructed signal and the ground-truth signal possess different physical units, the standard signal-to-noise ratio (SNR) is not an appropriate measure of similarity. Instead, we first compute the normalized cross-correlation to assess waveform fidelity. This metric offers a robust measure of similarity that remains invariant to differences in signal scale and DC offset. To further assess spectral fidelity, we compute the Median Vector Angle (MVA) error, following the methodology of Howard et al. Howard and Hirakawa (2023); Howard, Jones, and Hirakawa (2025). This metric quantifies spectral similarity while remaining robust to differences in the scale of the amplitude and the global phase. A lower MVA error indicates a higher correlation in the frequency domain. The calculation first determines the cosine similarity, $M(j)$, between the short-time Fourier transform of the reconstructed signal, $\mathcal{G}(j,\omega)$, and that of the ground-truth reference signal, $\mathcal{F}(j,\omega)$, for each short-time index $j$.
@@
Using the optimized ROI, we evaluated our method against two baselines: a passive event-based approach (Niwa et al. Niwa et al. (2023)) and a frame-based method (Davis et al. Davis et al. (2014)), both of which are foundational works in event- and frame-based vibration sensing. Our focus was on measuring vibrations on diffuse-reflective surfaces, which are common in real-world scenarios. Because our target exhibits diffuse rather than specular reflection, the method of Howard et al. Howard and Hirakawa (2023); Howard, Jones, and Hirakawa (2025), which relies on specular reflection, was not applicable and was therefore excluded from comparison. For the frame-based baseline (Davis et al.), we converted the event data into images by generating accumulated event frames, which is a standard technique in which events are integrated over time Gallego et al. (2022). To facilitate direct visual comparison in Figure 4, all reconstructed waveforms were mean-centered and normalized to the range [1,1][-1,1]. This preprocessing step that does not affect the quantitative evaluation, as both normalized cross-correlation and MVA error are invariant to such linear transformations.
@@
Signal Abe et al.
Niwa et al.
Ours
@@
Method Abe et al.
Niwa et al.
Ours

### v2

@@ Abstract
Event cameras are also particularly well suited for vibration measurement owing to their ability to capture fast dynamics. Approaches for vibration reconstruction using an event camera can broadly be categorized as active or passive capture. Active methods, which project an external light source onto a target, can realize high-sensitivity recordings. However, their use of lasers introduces limitations Yin et al. (2025); Cai et al. (2025). Although this approach is precise for single-point measurements, its reliance on capturing laser speckle patterns introduces practical limitations. For instance, the need to defocus the laser makes it difficult to record multiple vibration sources simultaneously, preventing the full utilization of the event camera’s high spatial resolution. Thus, we focus on passive methods, which record the target without using a strong light source. One approach to vibration reconstruction is interpreting the times of specular reflections in event data as signal level crossings and recovering the signal through short-time Fourier sparsity Howard and Hirakawa (2023); Howard et al. (2025). In our previous work, we also reconstructed vibrations by generating pseudo-frames from the event data and extracting phase information using Gabor filters Niwa et al. (2023). Although these passive methods can successfully reconstruct the principal frequency components of a vibration, they fail to accurately recover vibration amplitudes.
@@
Next, a clustering algorithm is applied to the subset of events within each colored interval. This process partitions the events into distinct clusters. For this task, we employ HDBSCAN (Hierarchical Density-Based Spatial Clustering of Applications with Noise) Campello et al. (2013). Unlike conventional density-based clustering such as DBSCAN Ester et al. (1996), which requires a manually tuned distance threshold $\varepsilon$, HDBSCAN adapts to varying event densities. This adaptability is crucial for robustly handling complex vibrations where frequency and amplitude changes alter the local event density. Figure 2(c) illustrates the outcome of this stage, showing the identified clusters after HDBSCAN has filtered out sparse events classified as noise. The interval and HDBSCAN parameters are detailed in the supplementary material.
@@
To evaluate the reconstruction accuracy, we employ two primary metrics. Because the reconstructed signal and the ground-truth signal possess different physical units, the standard signal-to-noise ratio (SNR) is not an appropriate measure of similarity. Instead, we first compute the normalized cross-correlation to assess waveform fidelity. This metric offers a robust measure of similarity that remains invariant to differences in signal scale and DC offset. To further assess spectral fidelity, we compute the Median Vector Angle (MVA) error, following the methodology of Howard et al. Howard and Hirakawa (2023); Howard et al. (2025). This metric quantifies spectral similarity while remaining robust to differences in the scale of the amplitude and the global phase. A lower MVA error indicates a higher correlation in the frequency domain. The calculation first determines the cosine similarity, $M(j)$, between the short-time Fourier transform of the reconstructed signal, $\mathcal{G}(j,\omega)$, and that of the ground-truth reference signal, $\mathcal{F}(j,\omega)$, for each short-time index $j$.
@@
Using the optimized ROI, we evaluated our method against two baselines: a passive event-based approach (Niwa et al. Niwa et al. (2023)) and a frame-based method (Davis et al. Davis et al. (2014)), both of which are foundational works in event- and frame-based vibration sensing. Our focus was on measuring vibrations on diffuse-reflective surfaces, which are common in real-world scenarios. Because our target exhibits diffuse rather than specular reflection, the method of Howard et al. Howard and Hirakawa (2023); Howard et al. (2025), which relies on specular reflection, was not applicable and was therefore excluded from comparison. For the frame-based baseline (Davis et al.), we converted the event data into images by generating accumulated event frames, which is a standard technique in which events are integrated over time Gallego et al. (2022). To facilitate direct visual comparison in Figure 4, all reconstructed waveforms were mean-centered and normalized to the range [1,1][-1,1]. This preprocessing step that does not affect the quantitative evaluation, as both normalized cross-correlation and MVA error are invariant to such linear transformations.
@@
Signal Abe et al. Niwa et al. (2023) Ours
@@
Method Abe et al. Niwa et al. (2023) Ours